OpenStudy (anonymous):

Find the radian measure of an angle of 110º

OpenStudy (anonymous):

110/360=x/2π x=2π*110/360 x=11π/18

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

welcome :)